A botanist observes within a wild population of sunflowers 34 tall plants, 12 medium-height plants, and 54 short plants. Sunflower height follows incomplete dominance, where tall (H) is dominant to short (h). We can calculate that the allele frequencies are:
Answer #1

Tall plants means TT=34

Medium height plants Tt= 12(T-6 and t-6)

Short plants or tt=54

Total individuals are 34+12+54= 100

In HW equilibrium, p is the dominant allele.

Here T is the dominant allele,and it's frequency is 34+6=40

Frequency 40/100=0.4

In HW, q is the recessive allele.

Here,t is the recessive allele and its frequency will be 54+6/100=0.6

Thus, p is 0.4 and q is 0.6

The population is in HW equilibrium because p+q=1